OpenStudy (anonymous):

Determine the answer to 3 + (−5) and explain the steps using a number line.

OpenStudy (danjs):

what are you having trouble with?

OpenStudy (anonymous):

i dont really understand how to find the answer

OpenStudy (danjs):

I think this wants to show you how a negative number moves left and positive to the right...so.. plus a negative number is the same as just subtraction 3 - 5

OpenStudy (danjs):

it is saying +(move right) by (-5) which itself moves left

OpenStudy (anonymous):

yea i know 3+(-5)= 3-5, but i still cant get the answer

OpenStudy (anonymous):

so what your saying is go right 3 times and then lest 5

OpenStudy (anonymous):

because i thought that was how u do it but i was confused

OpenStudy (horsegal244):

No its a negative

OpenStudy (danjs):

no just remember, a + (-b) = a - b plus a negative is subtraction

OpenStudy (anonymous):

so is the answer -2

OpenStudy (danjs):

if you add a negative number to a value on the numberline, you move left, same as if you just had subtracted it

OpenStudy (danjs):

the other one is minus a negative number, same as just addition a - (-b) = a + b

OpenStudy (anonymous):

i really need to get a good grade on this test. i did the number line and went from 0 to 3 then i went left 5 times and i got -2

OpenStudy (anonymous):

is that the right answer?

OpenStudy (danjs):

oh number line... forgot

OpenStudy (danjs):

3 + (−5) start at 3, move left 5 to -2

OpenStudy (anonymous):

so then i got the correct answer?

OpenStudy (danjs):



OpenStudy (anonymous):

i dont get that

OpenStudy (anonymous):

i started at 0 and went right till i got to the number 3. then i went left 5 times but started at 3 instead of 0 and i got -2. Is this right or wrong?

OpenStudy (danjs):



OpenStudy (danjs):

the upper one is just showing the distance between 3 and -2, not sure what you have to include on your graph

OpenStudy (anonymous):

just explain how to get the answer... or how u got the answer using the number line

OpenStudy (anonymous):

the explanation that i said earlier was it right or wrong

OpenStudy (danjs):

oh they want you to start at zero? then to show 3 - 5 = -2 ... start at zero, move +3, then move -5, end on -2

OpenStudy (anonymous):

thank you. i really needed to make sure i had the right answer

OpenStudy (anonymous):

can you help me with something else?

OpenStudy (danjs):

sure, i have a couple min

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Andrew is showing his work in simplifying −4.5 + 4.2 + 5.6 − 7.3. Identify any errors in his work or in his reasoning. Write feedback to Andrew explaining the mistake he made and how to correct his work. Step Work Reasoning −4.5 + 4.2 + 5.6 − 7.3 Original problem 1 −4.5 + 4.2 + 5.6 + (−7.3) Additive inverse 2 −4.5 + (−7.3) + 4.2 + 5.6 Commutative property 3 (−4.5 + (−7.3)) + (4.2 + 5.6) Associative property 4 −11.8 + 9.8 Simplify 5 −21.6 Simplify

OpenStudy (danjs):

k..

OpenStudy (anonymous):

in the original problem the 7.3 is positive, why in 1. he wrote it was negative

OpenStudy (danjs):

The 7.3 is negative in the original

OpenStudy (danjs):

the only thing i see is in the simplifying step

OpenStudy (danjs):

-11.8 + 9.8 = -2 not what he wrote -21.6